# Reasons to Reject? Aligning Language Models with Judgments

### v2

@@ 5.2.2 Training A Judgment Model
In the previous experiments, we show that CUT is effective in aligning LLMs with judgments annotated by humans or GPT4. However, both annotations can be expensive or infeasible. Therefore, we investigate the possibilities of developing an AI judge based on the current open-source LLMs.
Setup. we train AI judges with different amounts of judgment data $\{3000,5000\}$ collected in § 5.2.1. Then, we sample 1000 new instructions from Stanford Alpaca, obtain the corresponding responses from the target model (i.e., LLaMA2-chat), and label judgments with our AI judges. These new judgment triplets are used to align the target model.
Results. Figure 3 shows that AI judge-5000, trained with 5000 judgment data, is beneficial for aligning the target LLM, which leads to improvements of 3.02 and 4.17 points compared to LLaMA2-chat on TruthfulQA and AlpacaEval respectively. In contrast, AI Judge-3000, using a smaller training dataset, shows limited effectiveness. The comparison suggests that training a capable AI judge necessitates a moderate number of high-quality training instances. As a result, it is feasible to train AI judges to align the LLM. However, the quality of the AI judge remains a crucial factor in determining the success of this endeavor.

### v3

@@ 5.2.1 Training A Judgment Model
In the previous experiments, we show that CUT is effective in aligning LLMs with judgments annotated by humans or GPT4. However, human annotations can be very expensive. The use of GPT4 assumes that a very strong LLM already exists. Next, we investigate the possibilities of developing an AI judge based on the target LLM.
Setup. we train AI judges with different amounts of judgment data $\{3000,5000\}$ collected in § 5.2. Then, we sample 1000 new instructions from Stanford Alpaca, obtain the corresponding responses from the target model (i.e., LLaMA2-chat-13b), and label judgments with our AI judges. These new judgment triplets are used to align the target model.
Results. Figure 3 shows that AI judge-5000, trained with 5000 judgment data, is beneficial for aligning the target LLM, which leads to improvements of 3.02 and 4.17 points compared to LLaMA2-chat-13b on TruthfulQA and AlpacaEval respectively. In contrast, AI Judge-3000, using a smaller training dataset, shows limited effectiveness. The comparison suggests that training a capable AI judge necessitates a moderate number of high-quality training instances. As a result, it is feasible to train AI judges to align the LLM. However, the quality of the AI judge remains a crucial factor in determining the success of this endeavor.